OpenStudy (anonymous):

A South-African sharp nosed frog set a record for a triple jump by traveling a distance of 10.3 m. Suppose the frog starts from the origin of a coordinate system and lands at a point whose coordinate on the y-axis is equal to -6.10 m. What angle does the vector of displacement make with the negative y-axis? Calculate the x component of the frog.

OpenStudy (lgbasallote):

 the question is something like that?

OpenStudy (lgbasallote):

i think you can solve for the x coordinate using pythagorean theorem

OpenStudy (lgbasallote):

\[x = \sqrt{(10.3)^2 + (-6.10)^2}\]

OpenStudy (lgbasallote):

then...when you have the value of x... just do \[\theta = \tan^{-1}(\frac xy)\] does that help?

OpenStudy (lgbasallote):

@zipp you there?